ABC,. Inc just paid a dividend of $23. The dividends are expected to grow by 20% in Years 1 and 2. After that, the dividends are expected to grow by 3% each year. If the required rate of return is 18%, what is today's price of the stock?
Answer #1

Solution:

D0 = $23

D1= 23*1.20= $27.60

D2= $27.60*1.20= $33.12

Value after year 2= $33.12*1.03/(18% - 3%) = $227.424

Today's stock price = 27.60/1.18 + 33.12/(1.18) ^2 + $227.424/(1.18)^2= $210.51( approx)